OpenStudy (anonymous):

Use the properties of logarithms to solve the given equation. (Enter your answers as a comma-separated list.) log(2) (6x − x^2) = 3

OpenStudy (anonymous):

\[\log_{2} (6x-x ^{2}) =3\]

OpenStudy (turingtest):

Sorry, let me try again raise 2 to the power of both sides: \[6x-x^2=2^3=8\] \[x^2-6x+8=0\] (x-2)(x-4)=0 x={2,4}

OpenStudy (anonymous):

thanks